-2/3(6/5-7/10)=17/20
Answer by checkley77(12844)

-2/3(6/5-7/10)=17/20
-2/3(6*2-7)/10=17/20
-2/3(12-7)/10=17/20
-2/3*5/10=17/20
-10/30 DOES NOT =17/20
SOMETHING IS WRONG HERE!
CORRECT & RESUBMIT.